s have accounts payable and aceria stoc wsth dest and common esquity, so it s no a What is the b. What is the shere amount of total liabilities and quity that ap C What is the d. What is the cu that appears on the b Wharh companys total debt? total balance sheet? balance of current assets on the firms lance of current liabilities on the firms balance e. What is the amount of accounts payable and accruals on its balan heep Cc onsider this as a single line item on the firms balance sheet.) f. What is the firms ne g. Wha t working capital? t is the firms net operating working capital? at is the explanation for the difference in your answers to parts f and g INCOME STATEMENT Its Byron Books Inc. recently reported $13 million of EBIT was $20.8 million, and its tax rate was 35%. What was its internet rom

Chapter 3 Financial Statements, Cash Flow, and Taxes (Hint: Write out the headings for an income statement, and alues. Then $ 13 million of net income by (1 -T 0.65 to find the pretax income. The diference must be interest expense. Use this same procedure to fill in the known sv e between EBIT and taxable income complete similar problems.) .3 OME STATEMENT Patterson Brothers recently reported an EBITDA of 575 million and net income of $2.1 million. It had $2.0 million of interest expense, and its corporate tax rate was 30%. What was its charge for depreciation and amortization? 3- STATEMENT OF STOCKHOLDEaS SOTY In its most recent financial statements, Nessler Inc. reported $75 million of net income and $825 million of retained earnings. The previous retained earnings were $754 n tioa. ow much in dividends were paid to shareholders during the year? Assume thai al dividends declared were actually paid. MVA Harper Industries has $900 million of common equity on its balance sheet, its stock common shares are currently outstanding? 5 price is $80 per share; and its market value added (MVA) is $50 million. How many 6 MVA Over the years, Masterson Corporations stockholders have provided $34,000,000 of nital when they purchased new issues of stock and allowed management to retain sonm has 2,000,000 shares of common stock outstanding, ar Mnstersons management a
Answer #1
Ans 3-2
in million $
Net Income $13
Add:Income Tax (20-13) 7
EBT (13/(1-.35)) 20
Less: EBIT 20.8
Interest expenses 0.8
Interest expenses is $.8 million
ans 3-3
eBITDA $7.50
net Income 2.1
Interest expenses 2
Income tax expenses (3*.3) 0.9
working
EBT (2.1/.7) 3
Depreciation & amortization=
7.5-2.1-2-.9 $2.50 ANSWER
ans 3-4
Ending Retained Earning balance=Opening Ret Ear+net Income -dividend
825=784+75-D
Dividend=784+75-825 34 answer
ans 3--5
no. of commom stock
(MVA+Common equity)/market stock price
950/80 11.875
11.875*1000000 11875000 shares
If any doubt please comment